OpenStudy (anonymous):

solving equations with two variables algebraically 1: 2x-y=5 3x-2y=9

OpenStudy (anonymous):

x=1, y=-3

OpenStudy (anonymous):

(1) 2x-y=5 (2) 3x-2y=9 | (2)-2*(1) (1) 2x-y=5 (2) -x = -1 <=> x=1 x in (1): 2*1-y=5 solve for y: y=-3